# LoRA: Low-Rank Adaptation of Large Language Models

### v1

@@ Abstract
The dominant paradigm of natural language processing consists of large-scale pre-training on general domain data and adaptation to particular tasks or domains. As we pre-train larger models, conventional fine-tuning, which retrains all model parameters, becomes less feasible. Using GPT-3 175B as an example, deploying many independent instances of fine-tuned models, each with 175B parameters, is extremely expensive. We propose Low-Rank Adaptation, or LoRA, which freezes the pre-trained model weights and injects trainable rank decomposition matrices into each layer of the Transformer architecture, greatly reducing the number of trainable parameters for downstream tasks. For GPT-3, LoRA can reduce the number of trainable parameters by 10,000 times and the computation hardware requirement by 3 times compared to full fine-tuning. LoRA performs on-par or better than fine-tuning in model quality on both GPT-3 and GPT-2, despite having fewer trainable parameters, a higher training throughput, and no additional inference latency. We also provide an empirical investigation into rank-deficiency in language model adaptations, which sheds light on the efficacy of LoRA. We release our implementation in GPT-2 at https://github.com/microsoft/LoRA.
@@ 1 Introduction
Many applications in natural language processing rely on adapting one large scale, pre-trained language model to multiple downstream applications. Such adaptation is usually done via fine-tuning, which updates all the parameters of the pre-trained model. The major downside of fine-tuning is that it requires storing as many parameters as in the original model. As larger models are trained every few months, this changes from a mere “inconvenience” for GPT-2 [32] or BERT-large [9] to a critical deployment challenge for GPT-3 [6] with 175 billion trainable parameters. [^1] While GPT-3 175B achieves non-trivial performance with few-shot learning, fine-tuning boosts its performance significantly as shown in Appendix A.
@@
Many researchers sought to mitigate this by adapting only some parameters or learning external modules for new tasks. This way, we only need to load a small number of task-specific parameters to the pre-trained model for each task, which greatly boosts the deployment efficiency. However, in practice, existing techniques either introduce inference latency [15, 34] by extending model depth or reduce the model’s usable sequence length [21, 19, 14, 25]. More importantly, these prior attempts sometimes fail to match the fine-tuning baselines, posing a trade-off between efficiency and model quality.
@@
We take inspiration from [20, 1] which show that the learned over-parametrized models in fact reside on a low intrinsic dimension. We hypothesize that the update matrices in language model adaptation also have a low “intrinsic rank”, leading to our proposed Low-Rank Adaptation (LoRA) approach. LoRA allows us to train every dense layer in a neural network indirectly by injecting and optimizing rank decomposition matrices of the dense layer’s update instead, while keeping the original matrices frozen, as shown in Fig. 1. Using GPT-3 175B as an example, we show that a very low rank (i.e., r in Fig. 1 can be one or two) suffices even when the full rank (i.e., d) is as high as 12288, making LoRA both space- and compute-efficient.
LoRA possesses several key advantages as follows.
@@
  • A single pre-trained model can be shared and used to build many small LoRA modules for different tasks. We can keep the shared original model in VRAM and efficiently switch tasks by replacing the matrices $A$ and $B$ in Fig. 1, which significantly reduces the storage requirement and task-switching overhead.
@@
  • It makes training more efficient and lowers the hardware barrier to entry by 3 times, since we do not need to calculate the gradients or maintain the optimizer states for most model parameters when using adaptive optimizers. Instead, we only optimize the injected low-rank matrices, which have much fewer parameters.
@@
  • Its simple linear design allows us to merge the update matrices with the original weights during deployment, introducing no inference latency.
@@
  • LoRA is orthogonal to prior techniques and can be combined with many of them (such as prefix-tuning). We provide an example in Appendix D.
Terminologies
@@
We make frequent references to the Transformer architecture and use the conventional terminologies for its dimensions. We call its hidden size, or the size of its activations, $d_{model}$. We use $W_{q}$, $W_{k}$, $W_{v}$, and $W_{o}$ to refer to the query/key/value/output projection matrices in the self-attention module. $W$ or $W_{0}$ refers to a pre-trained weight matrix and $\Delta W$ its update during adaptation. We use $r$ to denote the rank of a LoRA module.
@@ 2 Problem Statement
While our proposal is agnostic to the training objective, we focus on language modeling as the primary use case. Below is a brief description of the language modeling problem and, in particular, the maximization of conditional probabilities given a task-specific prompt.
Suppose we are given a pre-trained autoregressive language model $p_{\Phi}(y|x)$ that is parametrized by $\Phi$. For instance, $p_{\Phi}(y|x)$ can be a generic multi-task learner such as GPT-2 [32] or GPT-3 [6] based on the Transformer architecture [36]. Now, consider adapting this pre-trained model to (possibly multiple) downstream conditional text generation tasks, such as summarization, machine reading comprehension (MRC), and natural language to SQL (NL2SQL). Each downstream task is represented by a training dataset of context-target pairs: $\mathcal{Z}=\{(x_{i},y_{i})\}_{i=1,..,N}$, where both $x_{i}$ and $y_{i}$ are sequences of tokens. For example, in NL2SQL, $x_{i}$ is a natural language query and $y_{i}$ its equivalent SQL command; for summarization, $x_{i}$ is the content of an article and $y_{i}$ its short summary.
During fine-tuning, the model is initialized with pre-trained parameters $\Phi_{0}$ and updated to $\Phi_{0}+\Delta\Phi$ by repeatedly following the gradient to maximize the conditional language modeling objective:
@@
One of the main drawbacks for full fine-tuning is that for each downstream task, we learn a different set of parameters $\Delta\Phi$ whose dimensions $|\Delta\Phi|$ equals $|\Phi_{0}|$. Thus, if the pre-trained model is large (such as GPT-3 with $|\Phi_{0}|\approx 175\text{~Billion}$), storing and deploying many independent instances of fine-tuned models can be challenging, if at all plausible.
In this paper, we adopt a parameter-efficient approach, where the task-specific parameter increment $\Delta\Phi=\Delta\Phi(\Theta)$ is further encoded by a much smaller-sized set of parameters $\Theta$ with $|\Theta|\ll|\Phi_{0}|$. The task of finding $\Delta\Phi$ thus becomes optimizing over $\Theta$:
@@
As we shall see in the subsequent sections, we propose to use a low-rank representation to encode $\Delta\Phi$ that is both computational and memory efficient. When the pre-trained model is GPT-3, the size of trainable parameters $|\Theta|$ can be as small as $0.01\%$ of $|\Phi_{0}|$.
@@ 3 Our Method
We describe the simple design of LoRA and its practical implications. The principles outlined here apply to any dense layers in deep learning models, though we only focus on certain weights in Transformers in our experiments for practical reasons.
Low-Rank Constraint on Update Matrices.  A typical neural network contains numerous dense layers that perform matrix multiplication. The weight matrices in these layers are allowed to have full-rank. When adapting to a specific task, however, Aghajanyan et al. [1] shows that the pre-trained language models have a low “instrisic dimension” and can still learn efficiently despite a low-dimensional reparametrization. Inspired by this observation, we wonder if the updates to the weights also have a low “intrinsic rank" when adapting to downstream tasks. For a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, we constrain its update by representing it with a low-rank decomposition $W_{0}+\Delta W=W_{0}+BA$, where $B\in\mathbb{R}^{d\times r},A\in\mathbb{R}^{r\times k}$, and the rank $r\ll min(d,k)$. During training, $W_{0}$ is frozen and does not receive gradient updates, while $A$ and $B$ contain trainable parameters. Note both $W_{0}$ and $\Delta W=BA$ are multiplied with the same input, and their respective output vectors are summed coordinate-wise. For $h=W_{0}x$, our modified forward pass yields:
@@
We illustrate our reparametrization in Fig. 1. We use a random Gaussian initialization for $A$ and zero for $B$, so $\Delta W=BA$ is zero at the beginning of training. We then scale $\Delta Wx$ by $\frac{1}{r}$ during training to keep the coordinates of $\Delta Wx$ roughly $\Theta(1)$ in $r$ after training [yang_feature_2021].
Weight Decay to Pre-trained Weights.  We note that weight decay behaves differently with LoRA than with full fine-tuning. Specifically, preforming the usual weight decay on $A$ and $B$ is similar to decaying back to the pre-trained weights, which has been studied as a potentially effective form of regularization against “catastrophic forgetting” [18, 16]. While extensive experiments isolating its effect is out-of-scope for this work, we believe that this, coupled with a constrained parameter space, might provide some regularization advantages. For example, see how LoRA with $r=d_{model}=1024$ outperforms full fine-tuning on GPT-2 Medium in Sec. 5.3 and Sec. G.2.
No Additional Inference Latency.  During deployment, we can explicitly compute $W=W_{0}+BA$ and perform inference as usual. When we need to switch to another downstream task, we can recover $W_{0}$ by subtracting $BA$ and then adding a different $B^{\prime}A^{\prime}$. This causes a minor increase in peak memory usage and adds a latency to model switching that does not exceed a single model forward pass. Critically, we do not introduce any additional latency during inference in return.
@@ 3.1 Applying LoRA to Transformer
In principle, we can apply LoRA to any subset of weight matrices in a neural network to reduce the number of trainable parameters. In the Transformer architecture, there are four weight matrices in the self-attention module ($W_{q},W_{k},W_{v},W_{o}$) and two in the MLP module. We treat $W_{q}$ (or $W_{k}$, $W_{v}$) as a single matrix of dimension $d_{model}\times d_{model}$, even though the output dimension is usually sliced into attention heads. We limit our study to only changing the attention weights for downstream tasks and freeze the MLP modules (so they are not trained in downstream tasks) since applying LoRA to the latter results in 4 times the number of trainable parameters given the same rank $r$. We further study the effect on adapting different types of weight matrices in a Transformer in Sec. 6.1.
@@
On the other hand, LoRA has its limitations. For example, it is not straightforward to batch inputs to different tasks with different $A$ and $B$ in a single forward pass, because we absorb $A$ and $B$ into $W$ to prevent additional inference latency.

### v2

@@ Abstract
An important paradigm of natural language processing consists of large-scale pre-training on general domain data and adaptation to particular tasks or domains. As we pre-train larger models, full fine-tuning, which retrains all model parameters, becomes less feasible. Using GPT-3 175B as an example – deploying independent instances of fine-tuned models, each with 175B parameters, is prohibitively expensive. We propose Low-Rank Adaptation, or LoRA, which freezes the pre-trained model weights and injects trainable rank decomposition matrices into each layer of the Transformer architecture, greatly reducing the number of trainable parameters for downstream tasks. Compared to GPT-3 175B fine-tuned with Adam, LoRA can reduce the number of trainable parameters by 10,000 times and the GPU memory requirement by 3 times. LoRA performs on-par or better than fine-tuning in model quality on RoBERTa, DeBERTa, GPT-2, and GPT-3, despite having fewer trainable parameters, a higher training throughput, and, unlike adapters, no additional inference latency. We also provide an empirical investigation into rank-deficiency in language model adaptation, which sheds light on the efficacy of LoRA. We release a package that facilitates the integration of LoRA with PyTorch models and provide our implementations and model checkpoints for RoBERTa, DeBERTa, and GPT-2 at https://github.com/microsoft/LoRA.
00footnotetext: Compared to V1, this draft includes better baselines, experiments on GLUE, and more on adapter latency.
@@ 1 Introduction
Many applications in natural language processing rely on adapting one large-scale, pre-trained language model to multiple downstream applications. Such adaptation is usually done via fine-tuning, which updates all the parameters of the pre-trained model. The major downside of fine-tuning is that the new model contains as many parameters as in the original model. As larger models are trained every few months, this changes from a mere “inconvenience” for GPT-2 (Radford et al., b) or RoBERTa large (Liu et al., 2019) to a critical deployment challenge for GPT-3 (Brown et al., 2020) with 175 billion trainable parameters. [^1] While GPT-3 175B achieves non-trivial performance with few-shot learning, fine-tuning boosts its performance significantly as shown in Appendix A.
@@
Many sought to mitigate this by adapting only some parameters or learning external modules for new tasks. This way, we only need to store and load a small number of task-specific parameters in addition to the pre-trained model for each task, greatly boosting the operational efficiency when deployed. However, existing techniques often introduce inference latency (Houlsby et al., 2019; Rebuffi et al., 2017) by extending model depth or reduce the model’s usable sequence length (Li & Liang, 2021; Lester et al., 2021; Hambardzumyan et al., 2020; Liu et al., 2021) (Section 3). More importantly, these method often fail to match the fine-tuning baselines, posing a trade-off between efficiency and model quality.
@@
We take inspiration from Li et al. (2018a); Aghajanyan et al. (2020) which show that the learned over-parametrized models in fact reside on a low intrinsic dimension. We hypothesize that the change in weights during model adaptation also has a low “intrinsic rank”, leading to our proposed Low-Rank Adaptation (LoRA) approach. LoRA allows us to train some dense layers in a neural network indirectly by optimizing rank decomposition matrices of the dense layers’ change during adaptation instead, while keeping the pre-trained weights frozen, as shown in Figure 1. Using GPT-3 175B as an example, we show that a very low rank (i.e., r in Figure 1 can be one or two) suffices even when the full rank (i.e., d) is as high as 12,288, making LoRA both storage- and compute-efficient.
LoRA possesses several key advantages.
@@
  • A pre-trained model can be shared and used to build many small LoRA modules for different tasks. We can freeze the shared model and efficiently switch tasks by replacing the matrices $A$ and $B$ in Figure 1, reducing the storage requirement and task-switching overhead significantly.
@@
  • LoRA makes training more efficient and lowers the hardware barrier to entry by up to 3 times when using adaptive optimizers since we do not need to calculate the gradients or maintain the optimizer states for most parameters. Instead, we only optimize the injected, much smaller low-rank matrices.
@@
  • Our simple linear design allows us to merge the trainable matrices with the frozen weights when deployed, introducing no inference latency compared to a fully fine-tuned model, by construction.
@@
  • LoRA is orthogonal to many prior methods and can be combined with many of them, such as prefix-tuning. We provide an example in Appendix E.
Terminologies and Conventions
@@
We make frequent references to the Transformer architecture and use the conventional terminologies for its dimensions. We call the input and output dimension size of a Transformer layer $d_{model}$. We use $W_{q}$, $W_{k}$, $W_{v}$, and $W_{o}$ to refer to the query/key/value/output projection matrices in the self-attention module. $W$ or $W_{0}$ refers to a pre-trained weight matrix and $\Delta W$ its accumulated gradient update during adaptation. We use $r$ to denote the rank of a LoRA module. We follow the conventions set out by (Vaswani et al., 2017; Brown et al., 2020) and use Adam (Loshchilov & Hutter, 2019; Kingma & Ba, 2017) for model optimization and use a Transformer MLP feedforward dimension $d_{ffn}=4\times d_{model}$.
@@ 2 Problem Statement
While our proposal is agnostic to training objective, we focus on language modeling as our motivating use case. Below is a brief description of the language modeling problem and, in particular, the maximization of conditional probabilities given a task-specific prompt.
Suppose we are given a pre-trained autoregressive language model $P_{\Phi}(y|x)$ parametrized by $\Phi$. For instance, $P_{\Phi}(y|x)$ can be a generic multi-task learner such as GPT (Radford et al., b; Brown et al., 2020) based on the Transformer architecture (Vaswani et al., 2017). Consider adapting this pre-trained model to downstream conditional text generation tasks, such as summarization, machine reading comprehension (MRC), and natural language to SQL (NL2SQL). Each downstream task is represented by a training dataset of context-target pairs: $\mathcal{Z}=\{(x_{i},y_{i})\}_{i=1,..,N}$, where both $x_{i}$ and $y_{i}$ are sequences of tokens. For example, in NL2SQL, $x_{i}$ is a natural language query and $y_{i}$ its corresponding SQL command; for summarization, $x_{i}$ is the content of an article and $y_{i}$ its summary.
During full fine-tuning, the model is initialized to pre-trained weights $\Phi_{0}$ and updated to $\Phi_{0}+\Delta\Phi$ by repeatedly following the gradient to maximize the conditional language modeling objective:
@@
One of the main drawbacks for full fine-tuning is that for each downstream task, we learn a different set of parameters $\Delta\Phi$ whose dimension $|\Delta\Phi|$ equals $|\Phi_{0}|$. Thus, if the pre-trained model is large (such as GPT-3 with $|\Phi_{0}|\approx 175\text{~Billion}$), storing and deploying many independent instances of fine-tuned models can be challenging, if at all feasible.
In this paper, we adopt a more parameter-efficient approach, where the task-specific parameter increment $\Delta\Phi=\Delta\Phi(\Theta)$ is further encoded by a much smaller-sized set of parameters $\Theta$ with $|\Theta|\ll|\Phi_{0}|$. The task of finding $\Delta\Phi$ thus becomes optimizing over $\Theta$:
@@
In the subsequent sections, we propose to use a low-rank representation to encode $\Delta\Phi$ that is both compute- and memory-efficient. When the pre-trained model is GPT-3 175B, the number of trainable parameters $|\Theta|$ can be as small as $0.01\%$ of $|\Phi_{0}|$.
3 Aren’t Existing Solutions Good Enough?
The problem we set out to tackle is by no means new. Since the inception of transfer learning, dozens of works have sought to make model adaptation more parameter- and compute-efficient. See Section 6 for a survey of some of the well-known works. Using language modeling as an example, there are two prominent strategies when it comes to efficient adaptations: adding adapter layers (Houlsby et al., 2019; Rebuffi et al., 2017; Pfeiffer et al., 2021; Rücklé et al., 2020) or optimizing some forms of the input layer activations (Li & Liang, 2021; Lester et al., 2021; Hambardzumyan et al., 2020; Liu et al., 2021). However, both strategies have their limitations, especially in a large-scale and latency-sensitive production scenario.
Adapter Layers Introduce Inference Latency
There are many variants of adapters. We focus on the original design by Houlsby et al. (2019) which has two adapter layers per Transformer block and a more recent one by  Lin et al. (2020) which has only one per block but with an additional LayerNorm (Ba et al., 2016). While one can reduce the overall latency by pruning layers or exploiting multi-task settings (Rücklé et al., 2020; Pfeiffer et al., 2021), there is no direct ways to bypass the extra compute in adapter layers. This seems like a non-issue since adapter layers are designed to have few parameters (sometimes <<1% of the original model) by having a small bottleneck dimension, which limits the FLOPs they can add. However, large neural networks rely on hardware parallelism to keep the latency low, and adapter layers have to be processed sequentially. This makes a difference in the online inference setting where the batch size is typically as small as one. In a generic scenario without model parallelism, such as running inference on GPT-2 (Radford et al., b) medium on a single GPU, we see a noticeable increase in latency when using adapters, even with a very small bottleneck dimension (Table 1).
Batch Size 32 16 1
Sequence Length 512 256 128
$|\Theta|$ 0.5M 11M 11M
[…]
Table 1: Infernece latency of a single forward pass in GPT-2 medium measured in milliseconds, averaged over 100 trials. We use an NVIDIA Quadro RTX8000. $|\Theta|$” denotes the number of trainable parameters in adapter layers. $\text{Adapter}^{\text{L}}$ and $\text{Adapter}^{\text{H}}$ are two variants of adapter tuning, which we describe in Section 5.1. The inference latency introduced by adapter layers can be significant in an online, short-sequence-length scenario. See the full study in Appendix B.
This problem gets worse when we need to shard the model as done in Shoeybi et al. (2020); Lepikhin et al. (2020), because the additional depth requires more synchronous GPU operations such as AllReduce and Broadcast, unless we store the adapter parameters redundantly many times.
Directly Optimizing the Prompt is Hard
The other direction, as exemplified by prefix tuning (Li & Liang, 2021), faces a different challenge. We observe that prefix tuning is difficult to optimize and that its performance changes non-monotonically in trainable parameters, confirming similar observations in the original paper. More fundamentally, reserving a part of the sequence length for adaptation necessarily reduces the sequence length available to process a downstream task, which we suspect makes tuning the prompt less performant compared to other methods. We defer the study on task performance to Section 5.
@@ 4 Our Method
We describe the simple design of LoRA and its practical benefits. The principles outlined here apply to any dense layers in deep learning models, though we only focus on certain weights in Transformer language models in our experiments as the motivating use case.
4.1 Low-Rank-Parametrized Update Matrices
A neural network contains many dense layers which perform matrix multiplication. The weight matrices in these layers typically have full-rank. When adapting to a specific task, Aghajanyan et al. (2020) shows that the pre-trained language models have a low “instrisic dimension” and can still learn efficiently despite a random projection to a smaller subspace. Inspired by this, we hypothesize the updates to the weights also have a low “intrinsic rank” during adaptation. For a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, we constrain its update by representing the latter with a low-rank decomposition $W_{0}+\Delta W=W_{0}+BA$, where $B\in\mathbb{R}^{d\times r},A\in\mathbb{R}^{r\times k}$, and the rank $r\ll\min(d,k)$. During training, $W_{0}$ is frozen and does not receive gradient updates, while $A$ and $B$ contain trainable parameters. Note both $W_{0}$ and $\Delta W=BA$ are multiplied with the same input, and their respective output vectors are summed coordinate-wise. For $h=W_{0}x$, our modified forward pass yields:
@@ 4.1 Low-Rank-Parametrized Update Matrices
We illustrate our reparametrization in Figure 1. We use a random Gaussian initialization for $A$ and zero for $B$, so $\Delta W=BA$ is zero at the beginning of training. We then scale $\Delta Wx$ by $\frac{\alpha}{r}$, where $\alpha$ is a constant in $r$. When optimizing with Adam, tuning $\alpha$ is roughly the same as tuning the learning rate if we scale the initialization appropriately. As a result, we simply set $\alpha$ to the first $r$ we try and do not tune it. This scaling helps to reduce the need to retune hyperparameters when we vary $r$ (Yang & Hu, 2021).
A Generalization of Full Fine-tuning.  A more general form of fine-tuning allows the training of a subset of the pre-trained parameters. LoRA takes a step further and does not require the accumulated gradient update to weight matrices to have full-rank during adaptation. This means that when applying LoRA to all weight matrices and training all biases [^2] They represent a negligible number of parameters compared to weights., we roughly recover the expressiveness of full fine-tuning by setting the LoRA rank $r$ to the rank of the pre-trained weight matrices. In other words, as we increase the number of trainable parameters  [^3] An inevitability when adapting to hard tasks., training LoRA roughly converges to training the original model, while adapter-based methods converges to an MLP and prefix-based methods to a model that cannot take long input sequences.
No Additional Inference Latency.  When deployed in production, we can explicitly compute and store $W=W_{0}+BA$ and perform inference as usual. Note that both $W_{0}$ and $BA$ are in $\mathbb{R}^{d\times k}$. When we need to switch to another downstream task, we can recover $W_{0}$ by subtracting $BA$ and then adding a different $B^{\prime}A^{\prime}$, a quick operation with very little memory overhead. Critically, this guarantees that we do not introduce any additional latency during inference compared to a fine-tuned model by construction.
@@ 4.2 Applying LoRA to Transformer
In principle, we can apply LoRA to any subset of weight matrices in a neural network to reduce the number of trainable parameters. In the Transformer architecture, there are four weight matrices in the self-attention module ($W_{q},W_{k},W_{v},W_{o}$) and two in the MLP module. We treat $W_{q}$ (or $W_{k}$, $W_{v}$) as a single matrix of dimension $d_{model}\times d_{model}$, even though the output dimension is usually sliced into attention heads. We limit our study to only adapting the attention weights for downstream tasks and freeze the MLP modules (so they are not trained in downstream tasks) both for simplicity and parameter-efficiency.We further study the effect on adapting different types of attention weight matrices in a Transformer in Section 7.1. We leave the empirical investigation of adapting the MLP layers, LayerNorm layers, and biases to a future work.
@@
LoRA also has its limitations. For example, it is not straightforward to batch inputs to different tasks with different $A$ and $B$ in a single forward pass, if one chooses to absorb $A$ and $B$ into $W$ to eliminate additional inference latency. Though it is possible to not merge the weights and dynamically choose the LoRA modules to use for samples in a batch for scenarios where latency is not critical.
5 Empirical Experiments
We evaluate the downstream task performance of LoRA on RoBERTa (Liu et al., 2019), DeBERTa (He et al., 2021), and GPT-2 (Radford et al., b), before scaling up to GPT-3 175B (Brown et al., 2020). Our experiments cover a wide range of tasks, from natural language understanding (NLU) to generation (NLG). Specifically, we evaluate on the GLUE (Wang et al., 2019) benchmark for RoBERTa and DeBERTa. We follow the setup of Li & Liang (2021) on GPT-2 for a direct comparison and add WikiSQL (Zhong et al., 2017) (NL to SQL queries) and SAMSum (Gliwa et al., 2019) (conversation summarization) for large-scale experiments on GPT-3. See Appendix C for more details on the datasets we use. We use NVIDIA Tesla V100 for all experiments.
5.1 Baselines
To compare with other baselines broadly, we replicate the setups used by prior work and reuse their reported numbers whenever possible. This, however, means that some baselines might only appear in certain experiments.
Fine-Tuning (FT) is a common approach for adaptation. During fine-tuning, the model is initialized to the pre-trained weights and biases, and all model parameters undergo gradient updates.A simple variant is to update only some layers while freezing others. We include one such baseline reported in prior work (Li & Liang, 2021) on GPT-2, which adapts just the last two layers ($\textbf{FT}^{\textbf{Top2}}$).
Bias-only or BitFit is a baseline where we only train the bias vectors while freezing everything else. Contemporarily, this baseline has also been studied by BitFit (Zaken et al., 2021).
Prefix-embedding tuning (PreEmbed) inserts special tokens among the input tokens. These special tokens have trainable word embeddings and are generally not in the model’s vocabulary. Where to place such tokens can have an impact on performance. We focus on “prefixing”, which prepends such tokens to the prompt, and “infixing”, which appends to the prompt; both are discussed in Li & Liang (2021). We use $l_{p}$ (resp. $l_{i}$) denote the number of prefix (resp. infix) tokens. The number of trainable parameters is $|\Theta|=d_{model}\times(l_{p}+l_{i})$.
Prefix-layer tuning (PreLayer) is an extension to prefix-embedding tuning. Instead of just learning the word embeddings (or equivalently, the activations after the embedding layer) for some special tokens, we learn the activations after every Transformer layer. The activations computed from previous layers are simply replaced by trainable ones. The resulting number of trainable parameters is $|\Theta|=L\times d_{model}\times(l_{p}+l_{i})$, where $L$ is the number of Transformer layers.
Adapter tuning as proposed in Houlsby et al. (2019) inserts adapter layers between the self-attention module (and the MLP module) and the subsequent residual connection. There are two fully connected layers with biases in an adapter layer with a nonlinearity in between. We call this original design $\textbf{Adapter}^{\textbf{H}}$. Recently, Lin et al. (2020) proposed a more efficient design with the adapter layer applied only after the MLP module and after a LayerNorm. We call it $\textbf{Adapter}^{\textbf{L}}$. This is very similar to another deign proposed in Pfeiffer et al. (2021), which we call $\textbf{Adapter}^{\textbf{P}}$. We also include another baseline call AdapterDrop (Rücklé et al., 2020) which drops some adapter layers for greater efficiency ($\textbf{Adapter}^{\textbf{D}}$). We cite numbers from prior works whenever possible to maximize the number of baselines we compare with; they are in rows with an asterisk (*) in the first column. In all cases, we have $|\Theta|=\hat{L}_{Adpt}\times(2\times d_{model}\times r+r+d_{model})+2\times\hat{L}_{LN}\times d_{model}$ where $\hat{L}_{Adpt}$ is the number of adapter layers and $\hat{L}_{LN}$ the number of trainable LayerNorms (e.g., in $\text{Adapter}^{\text{L}}$).
LoRA adds trainable pairs of rank decomposition matrices in parallel to existing weight matrices. As mentioned in Section 4.2, we only apply LoRA to $W_{q}$ and $W_{v}$ in most experiments for simplicity. The number of trainable parameters is determined by the rank $r$ and the shape of the original weights: $|\Theta|=2\times\hat{L}_{LoRA}\times d_{model}\times r$, where $\hat{L}_{LoRA}$ is the number of weight matrices we apply LoRA to.
5.2 RoBERTa base/large
Model & Method # Trainable
Parameters MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B Avg.
$\text{RoB}_{\text{base}}$ (FT)* 125.0M 87.6 94.8 90.2 63.6 92.8 91.9 78.7 91.2 86.4
$\text{RoB}_{\text{base}}$ (BitFit)* 0.1M 84.7 93.7 92.7 62.0 91.8 84.0 81.5 90.8 85.2
[…]
$\text{RoB}_{\text{large}}$ (FT)* 355.0M 90.2 96.4 90.9 68.0 94.7 92.2 86.6 92.4 88.9
[…]
$\text{DeB}_{\text{XXL}}$ (FT)* 1500.0M 91.8 97.2 92.0 72.0 96.0 92.7 93.9 92.9 91.1
[…]
Table 2: $\text{RoBERTa}_{\text{base}}$, $\text{RoBERTa}_{\text{large}}$, and $\text{DeBERTa}_{\text{XXL}}$ with different adaptation methods on the GLUE benchmark. We report the overall (matched and mismatched) accuracy for MNLI, Matthew’s correlation for CoLA, Pearson correlation for STS-B, and accuracy for other tasks. Higher is better for all metrics. * indicates numbers published in prior works. $\dagger$ indicates runs configured in a setup similar to Houlsby et al. (2019) for a fair comparison.
RoBERTa (Liu et al., 2019) optimized the pre-training recipe originally proposed in BERT (Devlin et al., 2019a) and boosted the latter’s task performance without introducing many more trainable parameters. While RoBERTa has been overtaken by much larger models on NLP leaderboards such as the GLUE benchmark (Wang et al., 2019) in recent years, it remains a competitive and popular pre-trained model for its size among practitioners. We take the pre-trained RoBERTa base (125M) and RoBERTa large (355M) from the HuggingFace Transformers library (Wolf et al., 2020) and evaluate the performance of different efficient adaptation approaches on tasks from the GLUE benchmark. We also replicate Houlsby et al. (2019) and Pfeiffer et al. (2021) according to their setup. To ensure a fair comparison, we make two crucial changes to how we evaluate LoRA when comparing with adapters. First, we use the same batch size for all tasks and use a sequence length of 128 to match the adapter baselines. Second, we initialize the model to the pre-trained model for MRPC, RTE, and STS-B, not a model already adapted to MNLI like the fine-tuning baseline. Runs following this more restricted setup from Houlsby et al. (2019) are labeled with $\dagger$. The result is presented in Table 2 (Top Three Sections). See Section D.1 for details on the hyperparameters used.
5.3 DeBERTa XXL
DeBERTa (He et al., 2021) is a more recent variant of BERT that is trained on a much larger scale and performs very competitively on benchmarks such as GLUE (Wang et al., 2019) and SuperGLUE (Wang et al., 2020). We evaluate if LoRA can still match the performance of a fully fine-tuned DeBERTa XXL (1.5B) on GLUE. The result is presented in Table 2 (Bottom Section). See Section D.2 for details on the hyperparameters used.
5.4 GPT-2 medium/large
Having shown that LoRA can be a competitive alternative to full fine-tuning on NLU, we hope to answer if LoRA still prevails on NLG models, such as GPT-2 medium and large (Radford et al., b). We keep our setup as close as possible to Li & Liang (2021) for a direct comparison. Due to space constraint, we only present our result on E2E NLG Challenge (Table 3) in this section. See Section F.1 for results on WebNLG (Gardent et al., 2017) and DART (Nan et al., 2020). We include a list of the hyperparameters used in Section D.3.
Model & Method # Trainable E2E NLG Challenge
Parameters BLEU NIST MET ROUGE-L CIDEr
GPT-2 M (FT)* 354.92M 68.2 8.62 46.2 71.0 2.47
GPT-2 M ($\text{Adapter}^{\text{L}}$)* 0.37M 66.3 8.41 45.0 69.8 2.40
GPT-2 M ($\text{Adapter}^{\text{L}}$)* 11.09M 68.9 8.71 46.1 71.3 2.47
[…]
GPT-2 M ($\text{FT}^{\text{Top2}}$)* 25.19M 68.1 8.59 46.0 70.8 2.41
GPT-2 M (PreLayer)* 0.35M 69.7 8.81 46.1 71.4 2.49
[…]
GPT-2 L (FT)* 774.03M 68.5 8.78 46.0 69.9 2.45
[…]
GPT-2 L (PreLayer)* 0.77M 70.3 8.85 46.2 71.7 2.47
[…]
Table 3: GPT-2 medium (M) and large (L) with different adaptation methods on the E2E NLG Challenge. For all metrics, higher is better. LoRA outperforms several baselines with comparable or fewer trainable parameters. Confidence intervals are shown for experiments we ran. * indicates numbers published in prior works.
Model&Method # Trainable WikiSQL MNLI-m SAMSum
Parameters Acc. (%) Acc. (%) R1/R2/RL
GPT-3 (FT) 175,255.8M 73.8 89.5 52.0/28.0/44.5
GPT-3 (BitFit) 14.2M 71.3 91.0 51.3/27.4/43.5
GPT-3 (PreEmbed) 3.2M 63.1 88.6 48.3/24.2/40.5
GPT-3 (PreLayer) 20.2M 70.1 89.5 50.8/27.3/43.5
GPT-3 ($\text{Adapter}^{\text{H}}$) 7.1M 71.9 89.8 53.0/28.9/44.8
GPT-3 ($\text{Adapter}^{\text{H}}$) 40.1M 73.2 91.5 53.2/29.0/45.1
GPT-3 (LoRA) 4.7M 73.4 91.7 53.8/29.8/45.9
GPT-3 (LoRA) 37.7M 74.0 91.6 53.4/29.2/45.1
Table 4: Performance of different adaptation methods on GPT-3 175B. We report the logical form validation accuracy on WikiSQL, validation accuracy on MultiNLI-matched, and Rouge-1/2/L on SAMSum. LoRA performs better than prior approaches, including full fine-tuning. The results on WikiSQL have a fluctuation around $\pm 0.5\%$, MNLI-m around $\pm 0.1\%$, and SAMSum around $\pm 0.2$$\pm 0.2$$\pm 0.1$ for the three metrics.
5.5 Scaling up to GPT-3 175B
As a final stress test for LoRA, we scale up to GPT-3 with 175 billion parameters. Due to the high training cost, we only report the typical standard deviation for a given task over random seeds, as opposed to providing one for every entry. See Section D.4 for details on the hyperparameters used.
As shown in Table 4, LoRA matches or exceeds the fine-tuning baseline on all three datasets. Note that not all methods benefit monotonically from having more trainable parameters, as shown in Figure 2. We observe a significant performance drop when we use more than 256 special tokens for prefix-embedding tuning or more than 32 special tokens for prefix-layer tuning. This corroborates similar observations in Li & Liang (2021). While a thorough investigation into this phenomenon is out-of-scope for this work, we suspect that having more special tokens causes the input distribution to shift further away from the pre-training data distribution. Separately, we investigate the performance of different adaptation approaches in the low-data regime in Section F.3.
Figure 2: GPT-3 175B validation accuracy vs. number of trainable parameters of several adaptation methods on WikiSQL and MNLI-matched. LoRA exhibits better scalability and task performance. See Section F.2 for more details on the plotted data points.